Are y=9x-14 and y=-1/10x-29 perpendicular

Answer by Cromlix(4381)

No, lines that are perpendicular
to one another have slopes that
multiply together to give -1
The slope of y = 9x - 14 is 9
The slope of y = -1/10x - 29 is -1/10
9 * -1/10 = -9/10 Not -1
If -1/10 had been -1/9
then 9 * -1/9 = -1.
Hope this helps.
:-)